Al’s Automotive started the year with total assets of $250,000 and total liabilities of $200,000. During the year the business recorded $225,000 in expenses, and had drawings of $35,000. Owner's equity at the end of the year was $110,000. What did the company record for revenue for the year?
Answer #1

Total assets=Total liabilities+Total equity

Total equity at beginning=(250,000-200,000)=$50,000

Ending equity=Beginning equity+Addition to retained earnings-Drawings

110,000=50,000+Addition to retained earnings-35000

Addition to retained earnings=(110,000+35000-50,000)

=$95000

Addition to retained earnings=Revenues-Expenses-Dividends

95000=Revenues-225,000-0

Revenues=(95000+225,000)

=$320,000